• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 16:00:36 on July22,2006

    Entry number 174835

    keyword=Day shift summary

    Shift Workers
    ---------------------------------------------------------------------
    SL: Yi Qiang
    TO: Xiaohui Zhan
    3rd: Roman Pomatsalyuk

    Shift Events
    ---------------------------------------------------------------------
    08:00 Still waiting for Roc3 ps to be fixed, jack is on his way
    09:00 Called Alex and Mike to come
    09:15 Tried to reset Q1 ps again with no success, so we paged John
    for help
    10:00 MCC called that our vacuum system had recovered
    10:30 Both Roc3 and Q1 ps problems are fixed
    11:00 MCC says they have some trouble to deliver high current beam and they still need some time to solve the problem
    15:00 Still waiting for beam
    16:00 End of shift with no beam. Good luck to following shift!